OpenStudy (anonymous):

Write the equation of the line that is parallel to the line 3x − y = −3 and passes through the point (4, −2).

OpenStudy (agent47):

parallel = same slope.

OpenStudy (agent47):

-y=-3x-3 y=3x+3 So you need slope 3

OpenStudy (agent47):

y-(-2)=3(x-4) Simplify, and you're good to go.

OpenStudy (anonymous):

IS THIS CORRECT y = 3x − 6

OpenStudy (paxpolaris):

no, \[\Large -2 \ne 3(4)-6\]

OpenStudy (paxpolaris):

y= 3x+c solve for c: \[\Large -2= 3(4)+c\]

OpenStudy (anonymous):

C=6

OpenStudy (paxpolaris):

\[\Large -2=12+c\] \[\implies \Large -2-12 = +c\]

OpenStudy (paxpolaris):

\[ \LARGE \implies-14=c\] y=3x-14